OpenStudy (miss.rose):

Luke mowed a lawn today to earn some extra money. He spent $2.10 on fuel and $4.45 on lunch. He charged $13.50 to mow the lawn. I will medal! Please help! @zoeeh127 Please help!

OpenStudy (tkhunny):

Add and subtract. Show your work.

OpenStudy (miss.rose):

So $2.10 + $4.45 = $6.55 The $13.50 - $6.55 = $6.95 Correct?

OpenStudy (anonymous):

Yes

OpenStudy (miss.rose):

Thanks! Can you help with: This morning, Elena wrote a check in the amount of $42.50. This brought her balance down to –$14.25. Suppose n represents Elena's balance before she wrote the check. Which equation can be used to represent this situation? n – 14.25 = –42.50 n – 14.25 = 42.50 n + 42.50 = –14.25 n – 42.50 = –14.25 . 1 points Question 2 A biologist is recording the loss of fish in a pond. He notes the number of fish, f, in the pond on June 1. On July 1 there were 63 fish in the pond, which is 52 fewer fish than were in the pond on June 1. Which equation represents this situation? f + 63 = –52 63 – f = 52 f + 63 = 52 f – 52 = 63